# Embed-Search-Align: DNA Sequence Alignment using Transformer models

### v3

@@ 3.1 Designing effective sequence representations
Here $i$ and $j$ serve to distinguish whether a read is aligned to a particular reference fragment, a positive sample $\{r_{j},\mathcal{F}_{j}\}$, or there is a mismatch (negative sample): $\{r_{j},\mathcal{F}_{i}\}$. Observe that these inequalities constitute the only requirements for the encoder. As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-K set) and alignment will succeed. Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment. This motivates using self-supervision (Hadsell et al., 2006; Chen et al., 2020a; Gao et al., 2021) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
@@ 3.2 Self-supervision and contrastive loss
A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  Chen et al. (2020a) and Gao et al. (2021): i.e. for a read $r$ aligned to reference fragment $\mathcal{F}_{j}$, the loss $l_{r}$ simultaneously minimizes the distance of $h(r)$ to $h(\mathcal{F}_{j})$ and maximizes the distance to a batch of random fragments of size $B-1$:
@@
Here $\tau$ is a tuneable temperature parameter. To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in  Chen et al. (2020b). Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (Reimers & Gurevych, 2019; Gao et al., 2021; Muennighoff et al., 2023), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
@@ 3.3 Encoder implementation
DNA-ESA uses a Transformer-encoder (Devlin et al., 2019; Vaswani et al., 2017), comprising $12$ heads and $6$-layers of encoder blocks. The size of vocabulary is $10,000$. Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps. The learning rate is annealed using one-cycle cosine annealing (Smith & Topin, 2019), dropout is set to $0.1$, and $\tau=0.05$. Reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}[800,2000]$ and read $|r_{i}|\sim\mathcal{U}[150,500]$. Shorter sequences were padded to equal the length of the longest sequence in a batch. The distance metric used is Cosine Similarity.
@@ 3.4 Search and retrieval
Figure 4: System Overview [B] - Inference on a New Read: A read, as per equation 1 and generated by ART (Huang et al., 2012), is encoded by $h$. This is then compared to reference fragment representations in the vector database. The nearest-K fragments in the embedding space are retrieved for each read, and the optimal alignment is determined using equation 6.
@@
Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments $\mathcal{F}:={\mathcal{F}_{1},\mathcal{F}_{2}},\dots$ to span $\mathcal{R}$. Note that the fragments overlap at least a read length; i.e. $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|\geq Q$ to guarantee that every read is fully contained within some fragment in the set. In our experiments with external read generators (Huang et al., 2012), $Q_{max}=250,|\mathcal{F}_{i}|=1250$. Each reference fragment is encoded using the trained DNA-ESA model, and the resulting sequence embeddings ($\in\mathbb{R}^{384}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database. Once populated with all the fragments, we are ready to perform the alignment.
@@
Diversity priors: While the top-$K$ retrieved fragments can be drawn from across the entire vector store (genome), contemporary recommendation systems that use the top-$K$ retrieval setup rank and re-rank top search results (Slate Optimization – see Zhu et al. (2007)) to ensure rich and diverse recommendations. Similarly, we apply a uniform prior wherein every retrieval step selects the top-$K$ per Chromosome.
Fine-Alignment: A standard SW distance library (Cock et al., 2009) is used to solve equation 3, which can be executed concurrently across the $K$-reference fragments. Let the optimal fragment be $\mathcal{F}^{*}$. The metadata for each vector includes (a) the raw $\mathcal{F}^{*}$ sequence; (b) the start position of $\mathcal{F}^{*}$ within the reference $\mathcal{R}$, $q_{\mathcal{F}^{*}|\mathcal{R}}$. Upon retrieval of a fragment and fine-alignment to find the fragment-level start index, $*q_{|\mathcal{F}^{*}}$, the global reference start position is obtained as:
@@ 4 Transformer-DNA baselines
This section outlines the setup for evaluating Transformer-DNA baselines, with their recall performance depicted in Fig. 2. We selected three architectures modeling nucleotide sequences: [NT] NucleotideTransformer ($\in\mathbb{R}^{1280}$(Dalla-Torre et al., 2023), [DB2] DNABERT-2 ($\in\mathbb{R}^{768}$(Ji et al., 2021), and [HD] HyenaDNA ($\in\mathbb{R}^{256}$(Nguyen et al., 2023). Each model employs mean- and max-pooling of token representations for sequence encoding ($2\times 3=6$ baselines total). Independent vector stores for each baseline encode fragments from the entire $3$gb genome. We sampled $40$K pure reads of varying lengths ($Q\sim\mathcal{U}[25,1000]$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig. 2. Overall, while baseline performance is modest, mean-pooling generally outperforms max-pooling, with DB2 (mean-pooled) and HD (max-pooled) as the most effective. These two baselines will be contrasted with DNA-ESA in Table 1.
5 Results and Discussion
@@
DNA-ESA convergence plots are presented in Figs. 4(a) and  4(b). Model checkpoints are available at OSF. In Fig. 1, representations of short $1,000-$length sequences sampled from sequential (in-order) and gene-specific locations in the reference are visualized in a reduced 2D-UMAP (McInnes et al., 2018). The representation space demonstrates desired properties suitable for successfully performing alignment: (a) Sequences sampled in order form a trajectory in the representation space: The loss function described in equation 5 encourages a pair of sequences close to one another to have a short distance between them in the representation space, and pairs further apart to have a larger distance. (b) Representations of sequences drawn from specific gene locations – despite not being close to one another – show gene-centric clustering: The DNA-ESA representation space partially acquires function-level separation as a byproduct of imposing local alignment constraints. Codebase is linked.
5.1 Sequence alignment of ART-simulated reads
@@
The results from Sec. 4 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment. In this section, DNA-ESA and the two best baselines – DB2, mean and HD, max – are evaluated using reads generated from an external read simulator (ART) – see Huang et al. (2012). ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.
@@
Simulator configurations: The different simulation configuration options and settings are listed: (A) Phred quality score $Q_{PH}\in\{[30,60],[60,90]\}$: the likelihood of errors in base-calls of a generated read; (B) Insertion rate $I\in\{0,10^{-2}\}$: the likelihood of adding a base to a random location in a read; (C) Deletion rate $D\in\{0,10^{-2}\}$: the likelihood of deleting a base in the read; (Others): Simulator system: MSv3 [MiSeq]; Read length: $250$.
@@
Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve equation 3, and for this, we need to compute the SW distance. For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1. After alignment, we get $q^{*}$ – see equation 6 – as the estimated location of a read in the genome. Let $\hat{q}^{*}$ be its true location. If $q^{*}=\hat{q}^{*}$, it is a perfect match and the recall is successful. In cases where there is a mutation in the first or last position in a read, the fine-alignment will return $q^{*}_{\mathcal{F}^{*}}$ offset by at most $2$ locations, resulting in $\hat{q}^{*}=q^{*}\pm 2$. Hence, the condition for an exact location match: $|q^{*}-\hat{q}^{*}|\leq 2$.
@@
Distance bound, $d_{SW}\in\{\text{None},1\%\}$: It is well known that short fragments frequently repeat in the genome (Li & Freudenberg, 2014) and $q^{*}$ can correspond to the position of the read in a different location than from where it was sampled. In this case, $q^{*}\neq\hat{q}^{*}$, but the SW distance is the minimum possible ($d_{SW}=0$ – see Footnote 2). Moreover, when reads have mutations, the reference sequence corresponding to the read is no longer a perfect match; i.e. $d_{SW}>0$. The best an alignment algorithm could do is to find an exact match for the read leading to an optimal alignment score for that read length ($-2Q$). We consider an alignment (with $Q=250$) to be successful if $d_{SW}<0.01$; i.e. a mismatch of at most $2$ bases.
DNA-ESA (ours)       DB2, mean HD, max
I D Top-$K$ 50 / Chr. $d_{SW}=1\%$ Best Top-K @ 50 / Chr. + $d_{SW}=1\%$
$Q_{PH}\in[60,90]$
0 0 95.2 ± 0.63 + 1.6 96.8 ± 0.52 37.4 ± 1.35 14.6 ± 1.01
0 0.01 95.8 ± 0.59 + 1.3 97.1 ± 0.50 37.4 ± 1.35 14.8 ± 1.01
0.01 0 95.1 ± 0.63 + 2.1 97.2 ± 0.49 35.9 ± 1.35 14.6 ± 1.01
0.01 0.01 95.8 ± 0.59 + 1.2 97.0 ± 0.51 34.9 ± 1.34 14.5 ± 1.01
$Q_{PH}\in[30,60]$
0 0 95.2 ± 0.63 + 1.4 96.6 ± 0.54 35.4 ± 1.34 13.7 ± 0.98
0 0.01 95.1 ± 0.64 + 2.0 97.1 ± 0.49 37.3 ± 1.35 14.2 ± 0.99
0.01 0 94.2 ± 0.68 + 2.1 96.3 ± 0.56 36.0 ± 1.35 13.8 ± 0.98
0.01 0.01 94.9 ± 0.64 + 1.5 96.4 ± 0.55 36.8 ± 1.35 13.5 ± 0.98
Table 1: Performance of DNA-ESA with respect to baselines: (with diversity priors – see Def. 4) The performance of DNA-ESA approeaches that of conventional algorithmic and far exceeds the performance of Transformer-DNA baselines, DB2, mean and HD, max (top performing baselines from Fig. 2). Both DNA-ESA and baselines utilize a dedicated vector store for the entire genome, maintaining a consistent search strategy. For details on $I,D,Q_{PH},K,d_{SW}$, refer to Recall/Simulator Configurations.
Performance: Table 1 contains the result of sweeps along these several parameters in addition to a direct comparison to DB2, mean and HD, max baselines, the best-performing baselines on pure reads in Sec. 4. We observe the following: (A) DNA-ESA demonstrates strong recall of $\sim 97\%$ across a variety of read generation and recall configurations described in Sec. 5.1; (B) Reads with less noise – high $Q_{PH}$, low $I,D$ – are more often correctly aligned; (C) The $d_{SW}$ bound adds $\sim 1.5-2\%$ in recall (to cover $5\%$ of misalignments) – a $\sim 20\%$ boost. This suggests that in the cases of misses, while finding the exact index match, the retrieved reference fragments are still high-quality retrievals that differ by at most $2$ bases; (D) While the precise method of (a) generating reads (ART settings, gapped, long vs. short, paired-end vs. single-end, etc.), (b) accounting for the variation in the reference (with or without mitochondrial DNA, single- or double haploid, etc.), (c) detecting a successful alignment (duplicates, distance criteria) vary considerably, DNA-ESA approaches the performance reported by mature algorithmic methods of StrobeAlign, BWA-Mem2, and Minimap (Sahlin, 2022; Vasimuddin et al., 2019; Li, 2018).

### v4

@@ 2.2 Designing effective sequence representations
Here $i$ and $j$ serve to distinguish whether a read is aligned to a particular reference fragment, a positive sample $\{r_{j},\mathcal{F}_{j}\}$, or there is a mismatch (negative sample): $\{r_{j},\mathcal{F}_{i}\}$. Observe that these inequalities constitute the only requirements for the encoder. As long as the neighborhood of $r_{j}$ in the representation space contains the representation for $\mathcal{F}_{j}$, it will be recovered in the nearest neighbors (top-$K$ set) and alignment will succeed. Equality is observed when $r_{j}$ is a repeat sequence matched equally well to more than one fragment. This motivates using self-supervision (hadsell_dimensionality_2006, ; chen_simple_2020, ; gao_simcse_2021, ) where we are only concerned about the relative distances between positive and negative (read, reference fragment) pairs.
@@ 2.3 Self-supervision and contrastive loss
A popular choice for sequence learning using self-supervision involves a contrastive loss setup described by  chen_simple_2020 and gao_simcse_2021 : i.e. for a read $r$ aligned to reference fragment $\mathcal{F}_{j}$, the loss $l_{r}$ simultaneously minimizes the distance of $h(r)$ to $h(\mathcal{F}_{j})$ and maximizes the distance to a batch of random fragments of size $B-1$:
@@
Here $\tau$ is a tuneable temperature parameter. To stabilize the training procedure and reach a non-trivial solution, the encoder applies different dropout masks to the reads and fragments similar to the method described in prior work simclr . Similar setups have been shown to work in written language applications, most notably in Sentence Transformers (reimers_sentence-bert_2019, ; gao_simcse_2021, ; muennighoff_mteb_2023, ), which continue to be a strong benchmark for several downstream tasks requiring pre-trained sequence embeddings.
@@ 2.4 Encoder implementation
DNA-ESA uses a Transformer-encoder (devlin_bert_2019, ; vaswani_attention_2017, ), comprising $12$ heads and $6$-layers of encoder blocks. The size of the vocabulary is $10,000$. Batch size $B$ is set to $16$ with gradient accumulation across $16$ steps. Generated numerical representations for each read, $h(r)\in\mathbb{R}^{1020}$. The learning rate is annealed using one-cycle cosine annealing (smith_super-convergence_2019, ), dropout is set to $0.1$, and $\tau=0.05$. During training, the reference fragment $|\mathcal{F}_{i}|\sim\mathcal{U}([800,2000])$ and read $|r_{i}|\sim\mathcal{U}([150,500])$ have variable lengths sampled from a uniform distribution. Here, $|x|$ denotes the length of $x$. To improve model performance on noisy reads, $1-5\%$ of bases are replaced with another random base in $40\%$ of the reads in a batch. Shorter sequences were padded to equal the length of the longest sequence in a batch. The distance metric used is Cosine Similarity.
@@ 2.5 Search and retrieval
Figure 4: System Overview [B] - Inference on a New Read: A read, as per (1) and generated by ART (huang_art_2012, ), is encoded by $h$. This is then compared to reference fragment representations in the vector database. The nearest-$K$ fragments in the embedding space are retrieved for each read, and the optimal alignment is determined using (7).
@@
Indexing: For a given reference genome $\mathcal{R}$, we construct a minimal set of reference fragments :=$\mathcal{F}:=\{\mathcal{F}_{1},\mathcal{F}_{2},\dots\}$ to span $\mathcal{R}$. Note that the fragments overlap at least a read length; i.e. $|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|\geq Q$ to guarantee that every read is fully contained within some fragment in the set. In our experiments with external read generators (huang_art_2012, ), $Q_{max}=250,|\mathcal{F}_{i}|=1250$. Each reference fragment is encoded using the trained DNA-ESA model, and the resulting sequence embeddings ($\in\mathbb{R}^{1020}$) – $3$M vectors for a reference of $3$B nucleotides – are inserted into a Pinecone database. Once populated with all the fragments, we are ready to perform the alignment.
@@
Diversity priors: While the top-$K$ retrieved fragments can be drawn from across the entire vector store (genome), contemporary recommendation systems that use the top-$K$ retrieval setup rank and re-rank top search results (Slate Optimization – see grasshopper ) to ensure rich and diverse recommendations. Similarly, we apply a uniform prior wherein every retrieval step selects the top-$K$ per chromosome.
Fine-Alignment: A standard SW distance library (biopython, ) is used to solve (4), which can be executed concurrently across the $K$-reference fragments. Let the optimal fragment be $\mathcal{F}^{*}$. The metadata for each vector includes (a) the raw $\mathcal{F}^{*}$ sequence; (b) the start position of $\mathcal{F}^{*}$ within the reference $\mathcal{R}$, $q_{\mathcal{F}^{*}|\mathcal{R}}$. Upon retrieval of a fragment and fine-alignment to find the fragment-level start index, $q_{|\mathcal{F}^{*}}$, the global reference start position is obtained as:
@@ 3 Transformer-DNA baselines
This section outlines the setup for evaluating Transformer-DNA baselines, with their recall performance depicted in Fig. 1. We selected three architectures modeling nucleotide sequences: [NT] NucleotideTransformer ($\in\mathbb{R}^{1280}$(nucleotide, ), [DB2] DNABERT-2 ($\in\mathbb{R}^{768}$(ji_dnabert_2021, ), and [HD] HyenaDNA ($\in\mathbb{R}^{256}$(nguyen_hyenadna_2023, ). Each model employs mean- and max-pooling of token representations for sequence encoding ($2\times 3=6$ baselines total). Independent vector stores for each baseline encode fragments from the entire $3$gb genome. We sampled $40$K pure reads of varying lengths ($Q\sim\mathcal{U}([25,1000])$) and assessed the average recall for top-$5$, top-$25$, and top-$50$ fragments, as shown in Fig. 1. Overall, while baseline performance is modest, mean-pooling generally outperforms max-pooling, with DB2 (mean-pooled) and HD (max-pooled) as the most effective. These two baselines will be contrasted with DNA-ESA in Table 1.
@@
DNA-ESA convergence plots are presented in Figs. 6(a) and  6(b). Model checkpoints are available at OSF. In Fig. 2, representations of short $1,000-$length sequences sampled from sequential (in-order) and gene-specific locations in the reference are visualized in a reduced 2D-UMAP (umap, ). The representation space demonstrates desired properties suitable for successfully performing alignment: (a) Sequences sampled in order form a trajectory in the representation space: The loss function described in (6) encourages a pair of sequences close to one another to have a short distance between them in the representation space, and pairs further apart to have a larger distance. (b) Representations of sequences drawn from specific gene locations – despite not being close to one another – show gene-centric clustering: The DNA-ESA representation space partially acquires function-level separation as a byproduct of imposing local alignment constraints. The codebase is linked.
4 Sequence alignment of ART-simulated reads
@@
The results from Sec. 3 demonstrate that even for pure reads, baseline models do not generate adequate representations to perform sequence alignment. In this section, DNA-ESA and the two best baselines – DB2, mean and HD, max – are evaluated using reads generated from an external read simulator (ART) – see huang_art_2012 . ART has served as a reliable benchmark for evaluating other contemporary alignment tools and provides controls to model mutations and variations common in reads generated by Illumina machines.
@@
Simulator configurations: The different simulation configuration options and settings are listed: (A) Phred quality score $Q_{PH}$ in one of three $\{[10,30],[30,60],[60,90]\}$ ranges: the likelihood of errors in base-calls of a generated read; (B) Insertion rate $I\in\{0,10^{-2}\}$: the likelihood of adding a base to a random location in a read; (C) Deletion rate $D\in\{0,10^{-2}\}$: the likelihood of deleting a base in the read; (Others): Simulator system: MSv3 [MiSeq]; Read length: $250$.
@@
Recall configurations: Once the top-K fragments have been retrieved, the first step is to solve (4), and for this, we need to compute the SW distance. For all presented results, the settings are: match_score = -2, mismatch_penalty = +1, open_gap_penalty = +0.5, continue_gap_penalty = +0.1. After alignment, we get $q^{*}$ – see (7) – as the estimated location of a read in the genome. Let $\hat{q}^{*}$ be its true location. If $q^{*}=\hat{q}^{*}$, it is a perfect match and the recall is successful. In cases where there is a mutation in the first or last position in a read, the fine-alignment will return $q^{*}_{\mathcal{F}^{*}}$ offset by at most $2$ locations, resulting in $\hat{q}^{*}=q^{*}\pm 2$. Hence, the condition for an exact location match: $|q^{*}-\hat{q}^{*}|\leq 2$.
@@
Distance bound, $d_{SW}\in\{1\%,2\%,5\%\}$: It is well known that short fragments frequently repeat in the genome and $q^{*}$ can correspond to the position of the read in a different location than from where it was sampled mappable ; reinventwillis . In this case, $q^{*}\neq\hat{q}^{*}$, but the SW distance is the minimum possible. Moreover, when reads have mutations, the reference sequence corresponding to the read is no longer a perfect match. We define a bound $d_{SW}$ for classifying whether a (read, retrieved fragment) pair is a successful alignment based on the SW distance between the pair ($v^{*}$ – see Eq. 4) and the optimal SW distance as a particular read length $Q$:
$d_{SW}=\frac{v^{*}-mQ}{(n-m)Q},$
where $m(=-2)$ is the match score and $n(=+1)$ is the mismatch penalty, hyperparameters in the computation of the SW distance. We consider an alignment (with $Q=250$) to be successful if the resulting SW distance between the read and the top returned fragment is within $d_{SW}$ of the optimal for that read length. A $d_{SW}=2\%$ (default) is equivalent to a mismatch of around $4$ bases in a read of length $250$.
Model Settings
ART
PacBio CCS
($\mu(v^{*}),\sigma(v^{*})$)
$I,D=0.01,Q=250$,
$d_{SW}=2\%$
Search [10,30] [30,60] [60,90]
Chr. 2, pbmm2
$d_{SW}=5\%$
Transformer Architectures HD (max) K=50 13.20 $\pm$ 1.81 17.01 $\pm$ 2.01 18.03 $\pm$ 2.05 -
DB2 (mean) 30.21 $\pm$ 2.44 39.19 $\pm$ 2.58 39.20 $\pm$ 2.58 -
DNA-ESA (ours) 98.40 $\pm$ 0.71 98.64 $\pm$ 0.67 98.60 $\pm$ 0.67 97.5 $\pm$ 0.82
K=75 98.80 $\pm$ 0.62 99.28 $\pm$ 0.52 98.88 $\pm$ 0.60
97.5 $\pm$ 0.82
(-498.0, 3.78)
Bowtie-2
- 99.80 $\pm$ 0.19
99.50$\pm$0.43
(-497.9, 3.78)
diff. - <1% <2%
Table 1: Performance of DNA-ESA with respect to baselines: The performance of DNA-ESA is compared to the DNA-BERT 2 and Hyena-DNA Transformer-based baseline models (with no additional finetuning) – the top performing baselines from Fig. 1. In addition, DNA-ESA is also compared with Bowtie-2 bowtie , a classical aligner. Comparisons are conducted across varying qualities (Phred score) of reads generated by ART huang_art_2012 . Additionally, an external PacBio CCS dataset of ($Q=250$ length-) reads from Chr. 2 of the Ashkenazim Trio - Son sample (as determined by pbmm2 minimap ) are aligned using both Bowtie-2 and DNA-ESA. All baseline models utilize a dedicated vector store. For details on $I,D,Q_{PH},K,d_{SW}$, refer to Recall/Simulator Configurations. Across models and ART-generated datasets, the performance of DNA-ESA supersedes other Transformer-based models and is comparable to Bowtie-2 to within $1\%$. With respect to reads from PacBio CCS, DNA-ESA performs with $2\%$ of Bowtie-2 after controlling for the intrinsic quality of retrieved fragments according to the SW distance. Mean ($\mu$) and standard deviation ($\sigma$) across the top SW distance values ($v^{*}$) of reads successfully aligned in the case of DNA-ESA and Bowtie-2 models are reported.
4.1 Performance
Table 1 reports the performance of DNA-ESA across several read generation configurations – we choose from one of three ranges of Phred score $Q_{PH}$, $\{[10,30],[30,60],[60,90]\}$$I,D=0.01$, $d_{SW}=2\%$ – see Sec. 4), in addition to a direct comparison to DB2, mean and HD, max baselines (without any further finetuning), the best-performing baselines on pure reads identified in Sec. 3. In addition, we also compare DNA-ESA to Bowtie-2 bowtie , a conventional algorithmic aligner.
Additionally, models are also evaluated on an external set of reads from PacBio CCN generated on the Ashkenazim Trio (Son) (retrieved from the Genome in a Bottle resource (GIAB) giab ). For (ii), the raw reads are $10$ kilobases long and for evaluation we consider random subset of $5000$ reads associated to Chr. 2333Reads specific to a chromosome are filtered using the pbmm2 (Minimap) package minimap . Reads input into the aligner are cropped randomly to a length of $250$ to satisfy the computational requirements of DNA-ESA ($|\mathcal{F}_{i}\bigcap\mathcal{F}_{i+1}|=250$).
We observe that: (A) DNA-ESA shows strong recall performance of $>99\%$ across a variety of read generation and recall configurations; (B) On cleaner reads, DNA-ESA and Bowtie-2 are within $<1\%$ of each other in recall while controlling for the quality of the retrieved alignments (as determined by the SW distance). The results indicate that this constitutes a new state-of-the-art model for Transformer-based sequence alignment.
Sweeping Top-K and the SW distance bound In Table 2, we report the performance of DNA-ESA across distance bounds $d_{SW}\in\{1\%,2\%,5\%\}$ and a number of recalled fragment settings $K\in\{25,50,75\}$. Distance bound $d_{SW}<5\%$ is equivalent to an acceptable mismatch of at most $\sim 8$ bases between the read (length $Q=250$) and recalled reference fragments. Recall rates are comparable with different $d_{SW}$ values suggesting that when a read is successfully aligned, it is usually aligned to an objectively best match. Decreasing the top-$K$ per chromosome from $50\rightarrow 25$ does not substantially worsen performance ($<1\%$), indicating that the optimal retrievals are usually the closest in the embedding space. Additional ablation studies are presented in the Appendix.
$Q_{PH}\in[30,60]$ $Q_{PH}\in[60,90]$
$I$ $D$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$ $d_{SW}<1\%$ $d_{SW}<2\%$ $d_{SW}<5\%$
Top-25 / chromosome
0.0 0.0 97±0.94 98.12±0.76 98.56±0.69 97.24±0.91 97.96±0.0.80 98.6±0.69
0.01 96.88±0.97 98±0.78 98.36±0.73 97.24±0.91 97.96±0.80 98.76±0.64
0.01 0.0 97.08±0.93 98.04±0.78 98.56±0.69 97±0.94 98.24±0.75 98.92±0.59
0.01 97.16±0.80 97.96±0.56 99.2±0.52 97.6±0.85 98.12±0.76 98.76±0.62
Top-50 / chromosome
0.0 0.0 97.52±0.86 99.04±0.57 99.08±0.57 98.2±0.75 98.88±0.59 99.16±0.52
0.01 97.48±0.87 99.16±0.52 99.08±0.55 97.8±0.82 98.76±0.62 99.04±0.57
0.01 0.0 98.24±0.75 98.64±0.67 99.12±0.55 98.28±0.75 98.92±0.59 99.28±0.49
0.01 97.72±0.83 98.64±0.67 99.36±0.46 97.84±0.92 98.6±0.66 99.0±0.57
Top-75/ chromosome
0.0 0.0 98.04±0.78 99.12±0.55 99.0±0.57 98.4±0.71 99.12±0.55 98.4±0.71
0.01 98.4±0.71 99.04±0.56 99.4±0.46 98.44±0.71 98.84±0.62 99.24±0.52
0.01 0.0 98.64±0.67 99.0±0.57 99.2±0.52 98.0±0.78 98.72±0.64 99.48±0.43
0.01 98.68±0.64 99.28±0.49 99.4±0.46 98.48±0.69 98.88±0.59 99.4±0.46
Table 2: Sequence alignment recall of DNA-ESA sweeping top-K and $d_{SW}$: The various parameters are described in Sec. 4. DNA-ESA presents a recall of $>99\%$ across several read configurations rivaling contemporary algorithmic models such as Bowtie. As expected, performance improves with larger search radius (top-$K$), higher quality reads ($Q_{PH}$) and large distance bound $d_{SW}$.